OpenStudy (anonymous):

3(y – 5) + 2 = 5 y = 4 y = 7 y = –4 y = 6

OpenStudy (anonymous):

y = 6 is u want it worked out Let me know

OpenStudy (anonymous):

3(y-5)=5-2 3(y-5)=3

OpenStudy (anonymous):

thnk u so much guys

OpenStudy (anonymous):

y-5=3/3 y-5=1

OpenStudy (anonymous):

y=6